### v1

# Simulations of Interacting Binary Systems - Pathways to Radio Bright GRB Progenitors

Abstract
Although the association of gamma-ray bursts (GRBs) with massive stellar death is on firm footing, the nature of the progenitor system and the key ingredients required for a massive star to produce a gamma-ray burst remain open questions. Here, we investigate the evolution of a massive star with a closely orbiting compact object companion using the stellar evolution code MESA. In particular, we examine how the companion influences the angular momentum and circumstellar environment near the end of the massive star life. We find that tidal effects can cause the compact object companion to significantly increase the angular momentum of the massive star, for orbital periods in the range of up to days. We model the density profile evolution of the massive star and discuss how tidal interactions may also lead to stripping of the outer stellar envelope in a way that can create an environment around the binary system that deviates from a typical wind density profile. We show how our results depend on the metallicity of the system, initial spin of the star, mass ratio, as well as accretion and dynamo prescriptions in the simulations. We conclude that these systems may be viable progenitors for radio-bright, long gamma-ray bursts.
1 Introduction
Gamma-ray bursts (GRBs) are the most luminous objects in the universe. The connection between so-called long GRBs (lGRBs; those with prompt gamma-ray emission lasting more than two seconds) and the deaths of massive stars is strongly supported, both theoretically and observationally, including the direct association of supernova events with many lGRBs (Woosley, 1993; MacFadyen & Woosley, 1999; Bloom et al., 2002; Hjorth et al., 2003; Woosley & Bloom, 2006; Woosley & Heger, 2006; Kumar et al., 2008a, b; Hjorth & Bloom, 2012; Lyman et al., 2017). While stars with high angular momentum and a stripped hydrogen envelope are necessary ingredients, the exact conditions required to produce an lGRB from a collapsing star (MacFadyen & Woosley, 1999; Yoon & Langer, 2005; Hirschi et al., 2005; Yoon et al., 2006; Woosley & Heger, 2006), and indeed the jet launching mechanism itself (e.g. Barkov & Komissarov, 2008; Komissarov & Barkov, 2009; Lloyd-Ronning et al., 2019a; King & Pringle, 2021), remain uncertain. It is an ongoing pursuit to determine and understand which stars make gamma-ray bursts, and why.
This mystery has been further deepened by the recent evidence (Lloyd-Ronning & Fryer, 2017; Lloyd-Ronning et al., 2019b; Chakraborty et al., 2023) that those GRBs with radio afterglows (so-called radio-bright GRBs) appear to have significantly longer-lasting prompt gamma-ray duration and higher isotropic equivalent energy than their radio-dark counterparts. It has been suggested (Lloyd-Ronning, 2022) that these radio bright GRBs may be a result of a massive star collapsing in a system with an influential compact object companion. Under the right conditions, the companion can spin up the massive star, providing it the angular momentum it needs to create a longer lasting jet and therefore longer-lasting prompt gamma-ray burst; this companion may also cause a more extended and dense circumburst environment, providing the surrounding gas needed for the GRB to shine brightly in the radio band.
In this paper, we test the hypothesis that long duration, radio-bright GRBs could result from a massive star collapsing in an interacting binary system with a closely orbiting compact object companion. In particular, we explore the influence of a black hole companion over a range of phase space varying orbital period, metallicity, initial spin of the star, mass ratio, accretion rate, magnetic fields, and numerical stopping conditions. We use the well-developed, open-source stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA) to perform this study. We present the parameter space in which this system will produce a highly-spinning massive star with a potentially dense extended circumburst environment - exactly the conditions needed to produce a radio-bright GRB.
Our paper is organized as follows: In §2, we describe the code and simulation set-up (with a convergence study discussed in our Appendix). In §3, we present our results and show the range of initial conditions that lead to a substantial spin up of the massive star in our systems. Additionally, we provide representative density profiles of the massive star and discuss the implications for the density profile of the circumbinary medium. In §4, we present discussion and caveats to our results. In §5, we summarize our main conclusions and discuss future work to support this effort.
2 Code and Simulation Set-up
We use the one-dimensional stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA) for our simulations (Paxton et al., 2011, 2013, 2015, 2018, 2019; Paxton, 2021; Jermyn et al., 2023)111We use MESA release r21.12.1.. MESA solves the coupled structure and composition equations simultaneously. The MESA equation of state (EOS) is a blend of the OPAL (Rogers & Nayfonov, 2002), SCVH (Saumon et al., 1995), FreeEOS (Irwin, 2004), HELM (Timmes & Swesty, 2000), PC (Potekhin & Chabrier, 2010), and Skye (Jermyn et al., 2021b) EOSes. Radiative opacities are primarily from OPAL (Iglesias & Rogers, 1993, 1996), with low-temperature data from Ferguson et al. (2005) and the high-temperature, and the Compton-scattering dominated regime by Poutanen (2017). Electron conduction opacities are from Cassisi et al. (2007) and Blouin et al. (2020). Nuclear reaction rates are from JINA REACLIB (Cyburt et al., 2010), NACRE (Angulo et al., 1999) and additional tabulated weak reaction rates Fuller et al. (1985); Oda et al. (1994); Langanke & Martínez-Pinedo (2000). Screening is included via the prescription of Chugunov et al. (2007). Thermal neutrino loss rates are from Itoh et al. (1996). Roche lobe radii in binary systems are computed using the fit of Eggleton (1983). Mass transfer rates in Roche lobe overflowing binary systems are determined following the prescription of Ritter (1988).
We employ the binary module in MESA because it enables simultaneous evolution of an interacting pair of stars undergoing transfer of mass and angular momentum (Paxton et al., 2015). We use the stellar rotation and tidal interaction treatment for the binary evolution from Marchant (2019) and Mink (2019). With these tools, we investigate the tidal regime where a black hole (BH) companion significantly influences the evolution of a massive star prior to collapse. Again, we are in particular searching for the parameter space for which the companion can spin up the massive star, and also produce tidal effects that may lead to a denser (relative to a single massive star), extended circumbinary medium into which the GRB jet will travel. We describe the details of our simulation set-up below.
Model name | [days] | [g cm-2 s-1] | [g cm-2 s-1] | ||||
MS25_BH15_T1.5_Z1.0e-4 | 25 | 15 | 1.5 | 18.86 | |||
MS25_BH15_T2.0_Z1.0e-4 | 25 | 15 | 2.0 | 22.85 | |||
MS25_BH15_T2.5_Z1.0e-4 | 25 | 15 | 2.5 | 26.51 | |||
MS25_BH15_T3.0_Z1.0e-4 | 25 | 15 | 3.0 | 29.94 | |||
MS25_BH15_T4.0_Z1.0e-4 | 25 | 15 | 4.0 | 36.27 | |||
MS25_BH15_T1.5_Z1.0e-2 | 25 | 15 | 1.5 | 18.86 | |||
MS25_BH15_T2.0_Z1.0e-2 | 25 | 15 | 2.0 | 22.85 | |||
MS25_BH15_T2.5_Z1.0e-2 | 25 | 15 | 2.5 | 26.51 | |||
MS25_BH15_T3.0_Z1.0e-2 | 25 | 15 | 3.0 | 29.94 | |||
MS25_BH15_T4.0_Z1.0e-2 | 25 | 15 | 4.0 | 36.27 | |||
MS25_BH10_T1.5_Z1.0e-4 | 25 | 10 | 1.5 | 18.04 | |||
MS25_BH10_T3.0_Z1.0e-4 | 25 | 10 | 3.0 | 28.63 | |||
MS25_BH10_T1.5_Z1.0e-2 | 25 | 10 | 1.5 | 18.04 | |||
MS25_BH10_T3.0_Z1.0e-2 | 25 | 10 | 3.0 | 28.63 |
2.1 MESA Input Physics
We evolve a Zero Aged Main Sequence (ZAMS) star with a point mass companion at different mass ratios, orbital periods222We consider orbital periods of the BH that lead to significant tidal interaction but do not consider common envelope or merger scenarios., metallicities, stellar rotations, and mass/angular momentum transfer schemes. Table 1 gives the choice of parameters for our simulations.
A short description of our initial set-up is as follows:
Mass ratio - We consider a 25 massive star with either a 10 or 15 BH companion as physically viable systems, created in situ or by dynamical capture. Systems with these mass ratios have been observed in X-ray binaries (Kelley et al., 1983; Levine et al., 1993, 2000; Tauris & van den Heuvel, 2006; Falanga et al., 2015). In future work, we perform population synthesis models (see, e.g., Cason et al. 2024) to place constraints on the numbers of these systems we expect in our universe. However, for now, we assume these systems exist in sufficient numbers and focus on their individual evolution.
Orbital Period - We consider orbital periods (and subsequently orbital separations) where the BH companion may tidally influence the massive star at a distance beyond the star’s radius outside of the Roche lobe radius of the massive star. Furthermore, at these separations, a Newtonian treatment to the gravitational interaction of the binary is sufficient because the general relativistic corrections to the orbit and tide are very small.
Metallicity - We consider two values, and , where metallicity is defined as the mass fraction elements heavier than helium relative to the total mass of the gas. We note that these span two orders of magnitude but remain lower than solar metallicity (), while it has been demonstrated that long GRBs tend to occur on average in lower metallicity environments (Fruchter et al., 1999; Le Floc’h et al., 2003; Fruchter et al., 2006; Levesque et al., 2010; Graham & Fruchter, 2013, 2017; Palla et al., 2019; Graham et al., 2019). Nonetheless, these values still provide the endpoints to a large dynamic range over which to explore dependencies on metallicity.
Spin Synchronization - We use the stellar rotation (in binary systems) settings in Marchant (2019), who model systems similar to what we are considering here. In particular, we turn on a spin synchronization flag which enables the tidal spin up of a massive star by its companion. We keep the default of no initial relaxation of the stellar rotation to the orbital period.
Massive star initial rotation - We consider a range of zero and non-zero initial rotations for the massive star. We initialize massive star non-zero rotation at a velocity given by the binary evolution equations in Hurley et al. (2000) and Hurley et al. (2002), used in the binary population synthesis code COSMIC Breivik et al. (2020). We take specific values from the COSMIC output discussed in Kenoly et al. (2023). Additionally, we consider higher velocity initial conditions for the massive star at half of the break-up velocity.
Mass transfer - For the simulations in this paper, the BH companion is primarily beyond the Roche lobe of the massive star and outside the regime where the choice of mass transfer scheme between the massive star and its companion is important. We employ the default setting for MESA (a Kolb-Ritter (Kolb & Ritter, 1990) mass transfer scheme) and further constrain the mass transfer to be lower than the Eddington limit. We note that there are many subtleties that come into play when choosing the mass transfer prescriptions. For example, Cehula & Pejcha (2023) find an intermediate solution between optically thin and optically thick regimes for Roche lobe overflow; their model finds a mass transfer rate about two times smaller than Kolb & Ritter (1990) and four times smaller than Marchant (2019) for a 30 , low metallicity star. However, again, because we are considering encounters where the initial separation distances are larger than the Roche lobe radius, these subtleties generally do not influence our main results.
We set the prescription for angular momentum loss through mass transfer according to the results of de Mink et al. (2013); Mink (2019); Marchant (2019), who carefully examine how angular momentum is lost and/or transferred between stars in a binary system, for different evolutionary stages, chemical composition, masses, separations, and magnetic field configurations.
Spruit-Taylor Dynamo - Through magnetic coupling between layers in the stellar interior, the Spruit-Taylor dynamo (Spruit, 2002) is a possible pathway to transport angular momentum within the massive star, potentially leading to a spin down. However, as we discuss below, its effect is more pronounced in stars less massive than those we consider here. Nonetheless, we have run simulations with this flag both on and off and find a negligible impact on our results (we discuss this further below).
Rotation in Convection Zone - In the case of rotation in the convection zone, angular momentum can be be transported through various processes internal to the star (Augustson & Mathis, 2019). These processes are controlled by the diffusion coefficients of Goldreich-Schubert-Fricke, Eddington-Sweet circulation, secular shear instability, and dynamical shear instability (see, e.g. Heger et al., 2000, for a description). We use the default MESA settings for rotation in the convection zone as well as the diffusion coefficients.
Stopping Condition - We identify the end point for our simulations as the stage prior to collapse when the carbon-oxygen core forms at 6-8 Myr with a corresponding rapid increase in central temperature and density within the star. We report the maximum angular momentum transfer to the star prior to this stage in the simulated evolution.
In the following, we present our set of binary parameters that lead to viable progenitors of radio bright lGRBs, particularly through significant stellar spin up over the massive star lifetime due to interaction with its companion. In our numerical simulations, we track the evolution of quantities that describe the dynamics of the binary system such as stellar and BH mass, orbital separation of the binary, and change in spin, orbital, and total angular momentum as well as stellar quantities through radial profiles in density and temperature. In the Appendix, we discuss convergence studies of representative models MS25_BH15_T3.0_Z1.0e-4 and MS25_BH15_T3.0_Z1.0e-2 of the tidal interaction. We show that these quantities do not change appreciably over stellar lifetime and model number at increasing mass resolution. Additionally, we identify the end point in our simulations prior to collapse in tracking the evolution of the spin up of the massive star with the formation of the carbon-oxygen core.
3 Results
3.1 Spin up of the Massive Star
For the binary parameters given in Table 1, our numerical simulations show a significant spin up of the massive star due to the tidal interaction with a BH, consistent with analytic estimates for long GRB binary progenitors (Lloyd-Ronning, 2022). In the following, we describe how these changes in spin angular momentum for the massive star depend on stellar metallicity, massive star to BH mass ratio, and initial stellar spin.
3.1.1 Dependence on Metallicity
In Fig. 1, we show the change in angular momentum of the massive star with initial orbital period and corresponding orbital separation for two stellar metallicities at (left panels) and (right panels). In the top row, we give the change in spin angular momentum or extent to which the star is spun up over the lifetime of the star, in units of initial total angular momentum . These interactions increase in spin up for shorter orbital periods, . This is consistent with a characterization of the tidal encounter through a strength parameter (Press & Teukolsky, 1977), for circular orbits at stellar-BH separation distance ,
for orbital timescale and stellar timescale . For a given star, encounters with smaller orbital periods will give rise to smaller or stronger tidal interactions. For our simulations in Table 1,
we consider binaries in the weak or partial disruption regime . Note that the limit corresponds to the tidal disruption limit where the dynamical time scales for the orbit and star are comparable. Our tidal calculations relating the spin up of massive star to orbital period for using the detailed stellar model described in Sec. 2,
agree with results of the spin up of polytropic stars in the weak tidal encounter regime for (Kochanek, 1992; Khokhlov et al., 1993; Cheng & Evans, 2013). Importantly, we find that for a given mass and initial orbital separation , massive star with lower metallicity (left panel) give rise to lower spin up than higher metallicity (right panel). This decrease in spin is consistent with the encounter strength parameter in Eq. 1, where stars with lower metallicity are more compact and give rise to weaker encounters (larger ) at fixed and orbital separation . In the bottom row, we show the evolution of orbital separation in units of . For all of these encounters, the massive star is tidally captured, with a decrease in . This is consistent with the orbital separation to tidal radius and mass ratio regime described in the parameter study simulating the interaction between stellar BH and main-sequence stars in Kremer et al. (2022).
The values of the final specific angular momentum of the massive star in cgs units range from to . This has important implications for the spin of the BH remnant when this massive star collapses, as well as the angular momentum in the disk that forms around the BH during/just after collapse. We consider this in our Discussion Section below.
3.1.2 Dependence on Mass Ratio
We examine the spin up of the massive star and change in orbital separation over the massive star lifetime for two mass ratios , shown in Fig. 2. For these simulations, we consider orbital periods of 1.5 and 3 days for a 25 massive star with a 10 and 15 BH, at stellar metallicities and . We find that the more extreme mass ratio encounters (with a 10 BH) lead to a larger spin up. This effect is more pronounced with smaller orbital periods, where stellar mass loss is greater due to the increase in encounter strength. This is consistent with results shown in de Mink et al. (2013) relating mass and angular momentum loss to mass ratio in interacting binaries.
However, we emphasize that for all of these encounters, the massive star is tidally distorted and there is a decrease in orbital separation of the system, leading to an increase in its spin angular momentum.
3.1.3 Dependence on Initial Spin of the Massive Star
We consider binary encounters where the massive star is initialized over a range of rotational velocities that include non-rotating , slowly rotating , and highly rotating stars. We use km/s from COSMIC output, as discussed in Kenoly et al. (2023). The break-up velocity is given by km/s for , respectively. Fig. 3 shows the stellar spin up and orbital separation at these varying initial rotation velocities, over the massive star lifetime. The evolution of and are similar for both cases of stellar metallicity. The spin for all three initial rotations reach close to the same value at the end of stellar lifetime. However, we note that when the massive star is initially rapidly rotating , this is the result of the massive star spinning down over its lifetime while increasing in orbital separation.
3.2 Stellar Density Profiles
In Figure 4, we show the density profile of the massive star at low and high metallicity ( and , respectively) for the case of an initial orbital period of three days. We expect that when tidal effects are not strong, earlier in the star’s life when the binary separation is still relatively large, the low metallicity star is more compact and the stellar density profiles are similar to single massive stars (black lines in Figure 4).
However, when the orbital separation decreases and tidal interactions become important, the companion can have an interesting effect on the density profile of the massive star (which in turn has important implications for the density profile of the circumbinary medium). As we see here, tidal interactions deform the star in such a way that it loses some mass and increases in radius with the likelihood, as discussed in Cheng & Evans (2013), for shock heating in the outer layers.
We expect that tidal interactions with an extended gas around a star may lead to ejection and therefore contribute to the density of the circumbinary environment. Provided that some of this gas settles into the direction into which the GRB jet (that is produced when the massive star collapses) propagates, this can lead to brighter radio emission since the radio flux scales as the square root of gas density333This is in the case of optically thin synchrotron emission. (Sari et al., 1998).
3.3 Additional Tests and Caveats
There are various mechanisms, such as the Spruit-Taylor dynamo, other magnetic coupling effects, or convection effects (Spruit, 2002; Heger et al., 2005), that can transport angular momentum out of the star throughout its lifetime. For example, Figure 3 of Heger et al. (2005) shows a pronounced decrease in angular momentum when magnetic fields are present, particularly for the pre-supernova stages of a star’s evolution. However, Denissenkov & Pinsonneault (2007) show that some of the assumptions required to invoke these effects are not always valid and, as a result, the loss of angular momentum can be significantly reduced. Potter et al. (2012) analyze the importance of the Spruit-Tayler dynamo in stars over a range of masses, and find that for stars above about 15 it becomes more difficult to sustain the dynamo, and as such the magnetic braking that normally occurs when this dynamo is active may less severe for more massive stars (particularly in the earlier stages of their evolution). We found that turning this flag on and off did not affect our results. We also investigated the effects of a rotating convection zone within the massive star. We began this investigation by setting the diffusion coefficients as prescribed in Heger et al. (2000) equal to 0. We find no change in the final angular momentum of the massive star at the end of its lifetime when this flag was turned on.
4 Discussion
We have investigated the properties of a massive star with a closely orbiting compact object companion and find a range of parameter space in which the massive star is spun up (i.e. ends its life with high angular momentum). We conjecture that this then is a plausible model for a highly spinning BH-disk system upon the massive star’s collapse, one that is capable of producing a viable GRB. We note that tidal synchronization is crucial for the massive star to spin up, especially if it begins with low initial angular momentum.
For a tidally locked binary system, the angular momentum of the massive star prior to collapse can be estimated by (Lloyd-Ronning, 2022):
for stellar mass , stellar radius , orbital separation , and where is the total mass of the binary system. The spin up of the massive star we find in our simulations in this work is consistent with this analytic result.
We note that there is additional loss of orbital angular momentum through gravitational radiation (Peters, 1964) for which we have not accounted. However, for the orbital separations we consider (Newtonian gravity regime), this is a negligible effect and does not significantly affect our results over the timescales we are considering (Mapelli, 2020; Beradze et al., 2020). Other effects (that we do not consider in this study) such as common envelope evolution and dynamical friction can also change the angular momentum loss timescale, significantly affecting the orbital evolution and, in turn, the final angular momentum of the massive star.
In connecting our progenitor system to a GRB - which results when the highly spinning massive star in our system collapses to a black hole - we have also not accounted for the potential loss of angular momentum upon collapse of the star. Here, we are making the assumption that a highly spinning massive star at the end of its life will produce a highly spinning BH-disk system. This is reasonably justified as shown in Jermyn et al. (2021a), who demonstrate that highly spinning massive stars in the ranges we consider end up with a roughly BH remnant, with a spin parameter , and with a long-lived disk (all elements to necessary to launch a relativistic GRB jet). Qin et al. (2018) use MESA to explore a range of conditions, in systems similar to the ones we consider here, where the spin of the second BH is high. They find that for very closely orbiting binaries, the spin of the second BH (i.e. the BH produced from the massive star in the systems we consider) is indeed near maximum (), particularly for low metallicity systems where mass loss from line driven winds is relatively small.
Therefore, under the reasonable assumption of a highly spinning massive star produces a highly spinning BH-disk system, this central engine with high angular momentum will then produce a longer lasting GRB jet. This, in turn, can manifest observationally as longer lasting GRB prompt emission, consistent with what is observed in the class of radio bright GRBs.
We suggest that in some cases (especially for higher metallicity stars with a “puffier” density profile or for more vigorous tidal interactions), a denser (relative to a single massive star) medium may be formed around the binary system to a larger radius and may potentially lead to a brighter radio afterglow, where the radio flux scales as roughly the square root of the circumbinary density (Sari et al., 1998).
The tidal interaction between the black hole and the massive star is the underlying mechanism of the spin up of our massive star, as well as the extended density profile that deviates from the typical profile of a single massive star. We note that Sciarini et al. (2024) have pointed out that tidal prescriptions in binary stellar evolution codes can sometimes either over or underestimate the strength of the tidal torque and so caution must be taken in over-interpreting results.
5 Conclusions
We have modeled the evolution of a massive star with a closely orbiting compact object companion. Our study is motivated by the interesting result that radio bright GRBs tend to have a longer prompt gamma-ray burst duration compared to radio dark GRBs. As such, we have explored systems that potentially lead to a massive star with more angular momentum (and therefore a longer lived accretion disk and jet which reflects the duration of the prompt emission). We have also discussed the dependence of the density profile of the massive star on binary interaction and stellar metallicity (which has implications for the density of the circumbinary medium).
Our main conclusions are as follows:
We find that massive stars with closely orbiting compact object companions are significantly spun up through tidal interactions, and end their life with high angular momentum, conducive to producing a highly spinning BH-disk system that could power a relatively long-lived GRB jet. The specific angular momentum of the star at the end of its life aligns with the expectation of the angular momentum needed for a roughly 10 star to launch a relativistic jet through the Blanford-Znajek process (Blandford & Znajek, 1977).
We find tidal effects leading to significant stellar spin up for orbital periods up to about 4 days, and over a range of mass ratios.
We find that the tidal spin up is more pronounced for higher metallicity stars for a given binary orbital period, due to the larger radius/extent of the star. However, the spin up is still significantly present in low metallicity systems.
We find that spin up depends strongly on the initial rotation of the massive star. Unsurprisingly, stars with more initial angular momentum are not spun up for the orbital periods we consider. In these cases, tidal syncing causes the massive star to spin down to synchronize with the orbital period of the system.
We present two sample density profiles of a massive star at two metallicities, illustrating the greater extent of the higher metallicity star. We discuss how this not only affects tidal interactions but also potentially impacts the density profile of the medium surrounding the binary.
The results of our work here inform future detailed studies into the physics of gamma-ray bursts. Specifically, we use the angular momentum of the massive star at the end of its life to inform our initial BH and disk angular momentum set-up in upcoming general relativistic magnetohydrodynamic simulations of the GRB jet produced by these systems We find specific angular momenta of the massive star in our closely orbiting binaries consistent with what others have employed in the literature for their BH-disk simulations (e.g. James et al., 2022).
Our work has defined a range of parameter space for specific binary systems that could be viable GRB progenitors, particularly for the sub-class of radio bright GRBs. Our results can be used to model populations of these systems and estimate their rates (A. Cason et al. in prep), validating whether they align with observed rates of long radio bright GRBs.
Acknowledgments
We thank the MESA community for helpful conversations and the many resources provided to help run MESA. This work was supported by the U. S. Department of Energy through Los Alamos National Laboratory (LANL). LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). Research presented was supported by the Laboratory Directed Research and Development program of LANL project number 20230115ER. We acknowledge LANL Institutional Computing HPC Resources under project w23extremex. Additional research presented in this article was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20210808PRD1. LA-UR-24-22983
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding author.

### v2

@@
 Abstract
-Although the association of gamma-ray bursts (GRBs) with massive stellar death is on firm footing, the nature of the progenitor system and the key ingredients required for a massive star to produce a gamma-ray burst remain open questions.
-Here, we investigate the evolution of a massive star with a closely orbiting compact object companion using the stellar evolution code MESA.
-In particular, we examine how the companion influences the angular momentum and circumstellar environment near the end of the massive star life.
-We find that tidal effects can cause the compact object companion to significantly increase the angular momentum of the massive star, for orbital periods in the range of up to days.
-We model the density profile evolution of the massive star and discuss how tidal interactions may also lead to stripping of the outer stellar envelope in a way that can create an environment around the binary system that deviates from a typical wind density profile.
-We show how our results depend on the metallicity of the system, initial spin of the star, mass ratio, as well as accretion and dynamo prescriptions in the simulations.
-We conclude that these systems may be viable progenitors for radio-bright, long gamma-ray bursts.
+Although the association of gamma-ray bursts with massive stellar death is on firm footing, the nature of the progenitor system and the key ingredients required for a massive star to produce a gamma-ray burst remain open questions.
+Here, we investigate the evolution of a massive star with a closely orbiting black hole using the stellar evolution code MESA.
+In particular, we examine how the companion influences the angular momentum of the massive star as it evolves over its lifetime under a range of orbital periods where the tidal effects are significant.
+We consider dependence on stellar metallicity, mass ratio, initial stellar spin, as well as accretion and dynamo prescriptions while tracking the mass loss and angular momentum in the simulations.
+We conclude that massive star-black hole binary systems with low metallicity massive stars under an initial orbital period range of 20-50 days may be viable GRB progenitors with post-collapse black hole masses of 5-10 and a black hole spin parameter .
 1 Introduction
@@
 Under the right conditions, the companion can spin up the massive star, providing it the angular momentum it needs to create a longer lasting jet and therefore longer-lasting prompt gamma-ray burst; this companion may also cause a more extended and dense circumburst environment, providing the surrounding gas needed for the GRB to shine brightly in the radio band.
-In this paper, we test the hypothesis that long duration, radio-bright GRBs could result from a massive star collapsing in an interacting binary system with a closely orbiting compact object companion.
-In particular, we explore the influence of a black hole companion over a range of phase space varying orbital period, metallicity, initial spin of the star, mass ratio, accretion rate, magnetic fields, and numerical stopping conditions.
+In this paper, we examine the influence of a black hole (BH) companion on a massive star over a range of phase space varying orbital period, metallicity, initial spin of the star, and mass ratio and discuss the implications this has for the final angular momentum of the star and the potential GRB it may produce.
+We identify a particular region of the binary parameter space where the massive star is subject to negligible mass loss from both wind and Roche-lobe overflow (RLOF) as the tidal effects significantly increase the spin angular momentum prior to stellar collapse at the end of its evolution.
 We use the well-developed, open-source stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA) to perform this study.
-We present the parameter space in which this system will produce a highly-spinning massive star with a potentially dense extended circumburst environment - exactly the conditions needed to produce a radio-bright GRB.
+We present the parameter space in which this system will produce a highly-spinning massive star.
+Previous work in stellar evolution modeling of these binary progenitors has provided many important details on whether or not the tidal interaction is significant enough to spin up a star under a wide range of parameter space.
+In a thorough investigation of mass and angular momentum loss, Detmers et al.
+(2008) find that the best candidates for Wolf-Rayet stars with a compact object companion are at low stellar metallicity, where stellar winds are weak and the orbital separations are stable without merger.
+Marchant et al.
+(2016) and Mandel & de Mink (2016) consider the interaction of massive stars in tight binaries under chemically homogenous evolution merger scenarios characterized by mixing and wind-driven mass loss due to rotation and tides.
+In addition there have been detailed studies looking at the influence of a low mass star on a massive star in a tight binary, exploring the conditions underwhich the low mass star can spin up the resultant black hole that the massive star produces upon collapse, but without merging (Moreno Méndez, 2022).
+In this work, we investigate a new region of parameter space in our choice of masses, stellar metallicities, and wide separations with orbital periods up to 100 days while exploring the limit of tidal interactions without significant mass loss due to RLOF.
 Our paper is organized as follows: In §2, we describe the code and simulation set-up (with a convergence study discussed in our Appendix).
 In §3, we present our results and show the range of initial conditions that lead to a substantial spin up of the massive star in our systems.
-Additionally, we provide representative density profiles of the massive star and discuss the implications for the density profile of the circumbinary medium.
+Additionally, we provide representative density profiles of the massive star and briefly discuss the implications for the density profile of the circumbinary medium.
 In §4, we present discussion and caveats to our results.
@@
 We use the stellar rotation and tidal interaction treatment for the binary evolution from Marchant (2019) and Mink (2019).
-With these tools, we investigate the tidal regime where a black hole (BH) companion significantly influences the evolution of a massive star prior to collapse.
-Again, we are in particular searching for the parameter space for which the companion can spin up the massive star, and also produce tidal effects that may lead to a denser (relative to a single massive star), extended circumbinary medium into which the GRB jet will travel.
+With these tools, we investigate the tidal regime where a BH companion significantly influences the evolution of a massive star prior to collapse.
 We describe the details of our simulation set-up below.
-Model name | [days] | [g cm-2 s-1] | [g cm-2 s-1] | ||||
-MS25_BH15_T1.5_Z1.0e-4 | 25 | 15 | 1.5 | 18.86 | |||
-MS25_BH15_T2.0_Z1.0e-4 | 25 | 15 | 2.0 | 22.85 | |||
-MS25_BH15_T2.5_Z1.0e-4 | 25 | 15 | 2.5 | 26.51 | |||
-MS25_BH15_T3.0_Z1.0e-4 | 25 | 15 | 3.0 | 29.94 | |||
-MS25_BH15_T4.0_Z1.0e-4 | 25 | 15 | 4.0 | 36.27 | |||
-MS25_BH15_T1.5_Z1.0e-2 | 25 | 15 | 1.5 | 18.86 | |||
-MS25_BH15_T2.0_Z1.0e-2 | 25 | 15 | 2.0 | 22.85 | |||
-MS25_BH15_T2.5_Z1.0e-2 | 25 | 15 | 2.5 | 26.51 | |||
-MS25_BH15_T3.0_Z1.0e-2 | 25 | 15 | 3.0 | 29.94 | |||
-MS25_BH15_T4.0_Z1.0e-2 | 25 | 15 | 4.0 | 36.27 | |||
-MS25_BH10_T1.5_Z1.0e-4 | 25 | 10 | 1.5 | 18.04 | |||
-MS25_BH10_T3.0_Z1.0e-4 | 25 | 10 | 3.0 | 28.63 | |||
-MS25_BH10_T1.5_Z1.0e-2 | 25 | 10 | 1.5 | 18.04 | |||
-MS25_BH10_T3.0_Z1.0e-2 | 25 | 10 | 3.0 | 28.63 |
+| [days] | ||||
+| 15.0 | 4.445 | 2.0 | 2.285 | |
+| 15.0 | 4.445 | 3.0 | 2.994 | |
+| 15.0 | 4.445 | 10.0 | 6.680 | |
+| 15.0 | 4.445 | 15.0 | 8.754 | |
+| 15.0 | 4.445 | 20.0 | 1.060 | |
+| 15.0 | 4.445 | 21.0 | 1.095 | |
+| 15.0 | 4.445 | 22.0 | 1.130 | |
+| 15.0 | 4.445 | 23.0 | 1.164 | |
+| 15.0 | 4.445 | 24.0 | 1.197 | |
+| 15.0 | 4.445 | 25.0 | 1.231 | |
+| 15.0 | 4.445 | 30.0 | 1.390 | |
+| 15.0 | 4.445 | 40.0 | 1.683 | |
+| 15.0 | 4.445 | 50.0 | 1.953 | |
+| 15.0 | 4.445 | 60.0 | 2.206 | |
+| 15.0 | 4.445 | 100.0 | 3.101 | |
+| 15.0 | 6.058 | 2.0 | 2.285 | |
+| 15.0 | 6.058 | 3.0 | 2.994 | |
+| 15.0 | 6.058 | 10.0 | 6.680 | |
+| 15.0 | 6.058 | 15.0 | 8.754 | |
+| 15.0 | 6.058 | 20.0 | 1.060 | |
+| 15.0 | 6.058 | 21.0 | 1.095 | |
+| 15.0 | 6.058 | 22.0 | 1.130 | |
+| 15.0 | 6.058 | 23.0 | 1.164 | |
+| 15.0 | 6.058 | 24.0 | 1.197 | |
+| 15.0 | 6.058 | 25.0 | 1.231 | |
+| 15.0 | 6.058 | 30.0 | 1.390 | |
+| 15.0 | 6.058 | 40.0 | 1.683 | |
+| 15.0 | 6.058 | 50.0 | 1.953 | |
+| 15.0 | 6.058 | 60.0 | 2.206 | |
+| 15.0 | 6.058 | 100.0 | 3.101 | |
+| 10.0 | 4.445 | 10.0 | 6.389 | |
+| 10.0 | 4.445 | 20.0 | 1.014 | |
+| 10.0 | 4.445 | 30.0 | 1.329 | |
+| 10.0 | 4.445 | 40.0 | 1.610 | |
+| 10.0 | 4.445 | 50.0 | 1.868 | |
+| 10.0 | 4.445 | 60.0 | 2.110 | |
+| 10.0 | 6.058 | 10.0 | 6.389 | |
+| 10.0 | 6.058 | 20.0 | 1.014 | |
+| 10.0 | 6.058 | 30.0 | 1.329 | |
+| 10.0 | 6.058 | 40.0 | 1.610 | |
+| 10.0 | 6.058 | 50.0 | 1.868 | |
+| 10.0 | 6.058 | 60.0 | 2.110 |
 2.1 MESA Input Physics
-We evolve a Zero Aged Main Sequence (ZAMS) star with a point mass companion at different mass ratios, orbital periods222We consider orbital periods of the BH that lead to significant tidal interaction but do not consider common envelope or merger scenarios., metallicities, stellar rotations, and mass/angular momentum transfer schemes.
+We evolve a Zero Aged Main Sequence (ZAMS) star in a circular orbit with a point mass BH at different mass ratios, orbital periods222We consider orbital periods of the BH that lead to significant tidal interaction but do not consider common envelope or merger scenarios., metallicities, stellar rotations, and mass/angular momentum transfer schemes.
 Table 1 gives the choice of parameters for our simulations.
@@
 Mass ratio - We consider a 25 massive star with either a 10 or 15 BH companion as physically viable systems, created in situ or by dynamical capture.
-Systems with these mass ratios have been observed in X-ray binaries (Kelley et al., 1983; Levine et al., 1993, 2000; Tauris & van den Heuvel, 2006; Falanga et al., 2015).
-In future work, we perform population synthesis models (see, e.g., Cason et al.
-2024) to place constraints on the numbers of these systems we expect in our universe.
+Systems like these, with these particular mass ratios, have been observed in X-ray binaries (Kelley et al., 1983; Levine et al., 1993, 2000; Tauris & van den Heuvel, 2006; Falanga et al., 2015).
+In future work, we perform population synthesis models (A.
+Cason et al.
+in prep) to place constraints on the numbers of these systems we expect in our universe.
 However, for now, we assume these systems exist in sufficient numbers and focus on their individual evolution.
-Orbital Period - We consider orbital periods (and subsequently orbital separations) where the BH companion may tidally influence the massive star at a distance beyond the star’s radius outside of the Roche lobe radius of the massive star.
+Orbital Period - We investigate orbital periods (and subsequently orbital separations) where the BH companion may tidally influence the massive star at a distance beyond the star’s radius outside of the Roche lobe radius of the massive star.
 Furthermore, at these separations, a Newtonian treatment to the gravitational interaction of the binary is sufficient because the general relativistic corrections to the orbit and tide are very small.
@@
 Nonetheless, these values still provide the endpoints to a large dynamic range over which to explore dependencies on metallicity.
-Spin Synchronization - We use the stellar rotation (in binary systems) settings in Marchant (2019), who model systems similar to what we are considering here.
-In particular, we turn on a spin synchronization flag which enables the tidal spin up of a massive star by its companion.
-We keep the default of no initial relaxation of the stellar rotation to the orbital period.
-Massive star initial rotation - We consider a range of zero and non-zero initial rotations for the massive star.
+Spin Synchronization - The binary module in MESA assumes that the star’s rotational axis is perpendicular to the orbital plane, enabling tidal interactions to contribute to the star’s spin-up through angular momentum transfer.
+MESA uses a modified scheme to solve for the star’s angular frequency as a function of time, following the formalism of Hut (1981) for a point-like mass companion while accounting for differential rotating stars (Paxton et al., 2015; Marchant, 2019).
+We do not initially relax the stellar rotation to the orbital period.
+Massive star initial rotation - We consider a range of stable, initial rotations for the massive star under the limit of stellar break-up velocity.
 We initialize massive star non-zero rotation at a velocity given by the binary evolution equations in Hurley et al.
 (2000) and Hurley et al.
-(2002), used in the binary population synthesis code COSMIC Breivik et al.
+(2002a), used in the binary population synthesis code COSMIC Breivik et al.
 (2020).
-We take specific values from the COSMIC output discussed in Kenoly et al.
-(2023).
-Additionally, we consider higher velocity initial conditions for the massive star at half of the break-up velocity.
-Mass transfer - For the simulations in this paper, the BH companion is primarily beyond the Roche lobe of the massive star and outside the regime where the choice of mass transfer scheme between the massive star and its companion is important.
-We employ the default setting for MESA (a Kolb-Ritter (Kolb & Ritter, 1990) mass transfer scheme) and further constrain the mass transfer to be lower than the Eddington limit.
+We initialize our simulations with zero velocity as a standard point in investigating the effects of spin.
+We also use two additional criteria for the initial velocity: The first ( km/s ) is set by the binary stellar evolution equations of Hurley et al.
+(2002b) and studied in Kenoly et al.
+(2023), who used the COSMIC population synthesis code (Breivik et al., 2020).
+The other is half of the critical velocity , defined as the velocity at which the star’s rotation rate will cause break up.
+Mass transfer - For the simulations in this paper, the BH companion is initially beyond the Roche lobe of the massive star and outside the regime where the choice of mass transfer scheme between the massive star and its companion is important.
+We employ a Kolb-Ritter (Kolb & Ritter, 1990; Paxton et al., 2015) mass transfer scheme, which accounts for stellar parameters such as finite atmosphere scale height and detailed structure of the outer layers.
+We further constrain the mass transfer to be lower than the Eddington limit.
 We note that there are many subtleties that come into play when choosing the mass transfer prescriptions.
-For example, Cehula & Pejcha (2023) find an intermediate solution between optically thin and optically thick regimes for Roche lobe overflow; their model finds a mass transfer rate about two times smaller than Kolb & Ritter (1990) and four times smaller than Marchant (2019) for a 30 , low metallicity star.
+For example, Cehula & Pejcha (2023) find an intermediate solution between optically thin and optically thick regimes for RLOF; their model finds a mass transfer rate about two times smaller than Kolb & Ritter (1990) and four times smaller than Marchant et al.
+(2016) for a 30 , low metallicity star.
 However, again, because we are considering encounters where the initial separation distances are larger than the Roche lobe radius, these subtleties generally do not influence our main results.
-We set the prescription for angular momentum loss through mass transfer according to the results of de Mink et al.
-(2013); Mink (2019); Marchant (2019), who carefully examine how angular momentum is lost and/or transferred between stars in a binary system, for different evolutionary stages, chemical composition, masses, separations, and magnetic field configurations.
-Spruit-Taylor Dynamo - Through magnetic coupling between layers in the stellar interior, the Spruit-Taylor dynamo (Spruit, 2002) is a possible pathway to transport angular momentum within the massive star, potentially leading to a spin down.
-However, as we discuss below, its effect is more pronounced in stars less massive than those we consider here.
+Furthermore, we note that in our simulations non-conservative mass transfer occurs such that mass from the massive star due to winds or RLOF does not necessarily end up accreted onto the black hole.
+In MESA, the corresponding loss in angular momentum is calculated from the net mass loss (Paxton et al., 2015; Marchant, 2019; Mink, 2019).
+Spruit-Tayler Dynamo - Through magnetic coupling between layers in the stellar interior, the Spruit-Tayler dynamo (Spruit, 2002; Fuller et al., 2019; Ma & Fuller, 2019) is a possible pathway to transport angular momentum within the massive star, potentially leading to a spin down.
+However, as we discuss below, its effect is more pronounced in stars less massive than those we consider here (see also the analytic arguments and discussion in Moreno Méndez, 2014, who show that for slowly spinning massive stars the slow-down from a Spruit-Taylor dynamo may be avoided).
 Nonetheless, we have run simulations with this flag both on and off and find a negligible impact on our results (we discuss this further below).
 Rotation in Convection Zone - In the case of rotation in the convection zone, angular momentum can be be transported through various processes internal to the star (Augustson & Mathis, 2019).
-These processes are controlled by the diffusion coefficients of Goldreich-Schubert-Fricke, Eddington-Sweet circulation, secular shear instability, and dynamical shear instability (see, e.g.
-Heger et al., 2000, for a description).
-We use the default MESA settings for rotation in the convection zone as well as the diffusion coefficients.
-Stopping Condition - We identify the end point for our simulations as the stage prior to collapse when the carbon-oxygen core forms at 6-8 Myr with a corresponding rapid increase in central temperature and density within the star.
-We report the maximum angular momentum transfer to the star prior to this stage in the simulated evolution.
-In the following, we present our set of binary parameters that lead to viable progenitors of radio bright lGRBs, particularly through significant stellar spin up over the massive star lifetime due to interaction with its companion.
-In our numerical simulations, we track the evolution of quantities that describe the dynamics of the binary system such as stellar and BH mass, orbital separation of the binary, and change in spin, orbital, and total angular momentum as well as stellar quantities through radial profiles in density and temperature.
-In the Appendix, we discuss convergence studies of representative models MS25_BH15_T3.0_Z1.0e-4 and MS25_BH15_T3.0_Z1.0e-2 of the tidal interaction.
-We show that these quantities do not change appreciably over stellar lifetime and model number at increasing mass resolution.
-Additionally, we identify the end point in our simulations prior to collapse in tracking the evolution of the spin up of the massive star with the formation of the carbon-oxygen core.
+These processes are controlled by the diffusion coefficients of Goldreich-Schubert-Fricke (GSF), Eddington-Sweet (ES) circulation, secular shear instability (SSI), and dynamical shear instability(DSI).
+The GSF (Goldreich & Schubert, 1967; Fricke, 1968) coefficient quantifies how angular momentum is transported in the interiors of stars.
+ES occurs when there are imbalances of pressure and temperature inside the interior of the star.
+SSI is a long-term instability in rotating stars where thermal diffusion weakens stratification, allowing shear-driven turbulence to redistribute angular momentum and mix elements.
+DSI occurs in rotating stars when shear forces exceed stabilizing buoyancy, triggering turbulence and rapid mixing on a dynamical timescale (Heger et al., 2000).
+Termination condition - We identify the end point for our simulations as the stage prior to collapse when the carbon-oxygen core forms at 6-8 Myr with a corresponding rapid increase in central temperature and density within the star.
+We also track the evolution of the helium core and central mass ratio of helium-4 and note that our stopping criteria for the simulations is consistent with this end point.
+We report the final angular momentum transfer to the star at this stage in the simulated evolution.
+In the following, we present our set of binary parameters that could lead to viable progenitors of radio bright lGRBs, particularly through significant stellar spin up over the massive star lifetime due to interaction with its companion.
+In our numerical simulations, we track the evolution of quantities that describe the dynamics of the binary system such as stellar and BH mass, orbital separation of the binary, and change in spin, orbital, and total angular momentum as well as stellar quantities such as mass loss due to wind and RLOF, stellar radius, RLOF radius, and rotation.
+We report the spin angular momentum and total mass loss for massive stars at the final stage of evolution indicated by our stopping criteria.
+In the Appendix, we discuss convergence studies of representative models of the binary interaction between a BH and massive star at an initial orbital period of three days to test our schemes where significant mass loss occurs.
+We show that our quantities of interest do not change appreciably over stellar lifetime and model number at increasing mass resolution.
 3 Results
-3.1 Spin up of the Massive Star
-For the binary parameters given in Table 1, our numerical simulations show a significant spin up of the massive star due to the tidal interaction with a BH, consistent with analytic estimates for long GRB binary progenitors (Lloyd-Ronning, 2022).
-In the following, we describe how these changes in spin angular momentum for the massive star depend on stellar metallicity, massive star to BH mass ratio, and initial stellar spin.
-3.1.1 Dependence on Metallicity
+For the binary parameters given in Tab.
+1, our numerical simulations show a particular region with significant spin up of the massive star due to the tidal interaction with a BH, consistent with analytic estimates for long GRB binary progenitors (Lloyd-Ronning, 2022).
 In Fig.
-1, we show the change in angular momentum of the massive star with initial orbital period and corresponding orbital separation for two stellar metallicities at (left panels) and (right panels).
-In the top row, we give the change in spin angular momentum or extent to which the star is spun up over the lifetime of the star, in units of initial total angular momentum .
-These interactions increase in spin up for shorter orbital periods, .
-This is consistent with a characterization of the tidal encounter through a strength parameter (Press & Teukolsky, 1977), for circular orbits at stellar-BH separation distance ,
-for orbital timescale and stellar timescale . For a given star, encounters with smaller orbital periods will give rise to smaller or stronger tidal interactions. For our simulations in Table 1,
-we consider binaries in the weak or partial disruption regime .
-Note that the limit corresponds to the tidal disruption limit where the dynamical time scales for the orbit and star are comparable.
-Our tidal calculations relating the spin up of massive star to orbital period for using the detailed stellar model described in Sec.
-2,
-agree with results of the spin up of polytropic stars in the weak tidal encounter regime for (Kochanek, 1992; Khokhlov et al., 1993; Cheng & Evans, 2013).
-Importantly, we find that for a given mass and initial orbital separation , massive star with lower metallicity (left panel) give rise to lower spin up than higher metallicity (right panel).
-This decrease in spin is consistent with the encounter strength parameter in Eq.
-1, where stars with lower metallicity are more compact and give rise to weaker encounters (larger ) at fixed and orbital separation .
-In the bottom row, we show the evolution of orbital separation in units of .
-For all of these encounters, the massive star is tidally captured, with a decrease in .
-This is consistent with the orbital separation to tidal radius and mass ratio regime described in the parameter study simulating the interaction between stellar BH and main-sequence stars in Kremer et al.
+1, we show the dependence of final stellar spin angular momentum and mass loss on initial orbital period with models given in Tab.
+1 for initially non-rotating stars.
+In Fig.
+2 and Fig.
+3, we provide an analysis of these models at initially three day and twenty-one day orbital periods and show the evolution of stellar radius, RLOF radius, mass loss due to stellar wind and RLOF, spin and orbital angular momentum, and our termination conditions in tracking the formation of the helium and carbon-oxygen core mass and central helium-4 mass fraction.
+In the following, we describe our results in detail with regards to the identification of the parameter space of viable progenitor binary systems, mass loss in late stage evolution due to RLOF, angular momentum, termination conditions, metallicity, mass-ratio, rotation, and additional tests and caveats.
+3.1 Viable progenitor binary systems with significant stellar spin-up and negligible mass loss
+We identify a region of parameter space where the tidal interaction in the binary leads to significant stellar spin-up such that after the massive star collapses, its remnant BH-disk system may retain enough angular momentum to launch a jet via Blandford-Znajek process (Blandford & Znajek, 1977).
+For a viable GRB from such a progenitor binary system, we estimate the lower limit for the final spin angular momentum as gcm-2s-1 using a BH mass of spin of 0.5.
+For a BH mass of , the lower limit is smaller by a factor of 4.
+In the first row of Fig.
+1, we show a region of parameter space in initial orbital period , stellar to BH mass ratios, and stellar metallicity where the final stellar spin-up is above the lower limit of a viable GRB progenitor with BH mass of (solid line) and (dotted line).
+We identify 25, low metallicity () massive stars interacting with both 15 and 10 BH for an initial orbital period of 20-50 days as viable progenitor binary systems.
+For this particular region of binary parameter space, we note negligible final total mass loss on the order of a factor of of the initial mass of the star.
+Outside of this region, non-viable binaries are characterized by significant final mass loss close to a factor of unity or at an orbital separation where the tidal interaction is negligible ( 60).
+Additionally, we note very slight decreases in orbital separation at the level of for simulations with negligible mass loss including the viable progenitor binary systems.
+This is consistent with Detmers et al.
+(2008), where it is shown that, while the spin-up process itself may work, significant mass loss due to winds and RLOF negate the spin-up effect through the corresponding loss in stellar angular momentum.
+Additionally, our results are comparable to the orbital separation to tidal radius and mass ratio regime described in the parameter study simulating the interaction between stellar BH and main-sequence stars in Kremer et al.
 (2022).
-The values of the final specific angular momentum of the massive star in cgs units range from to .
+3.2 Roche-lobe overflow in late stage stellar evolution
+For all simulations in Tab.
+1, the radius of the massive star increases significantly in the late-stages of stellar evolution until termination at Myr.
+While the RLOF radius remains relatively constant, there is an increase in the ratio between stellar to RLOF radius, leading to significant mass loss.
+We show that the primary drivers of mass loss are due to stellar winds and RLOF in simulations between a massive star and a 15 BH with an initial orbital period of three days in Fig.
+2 and twenty-one days in Fig.
+3.
+In the first row of both figures, we compare the ratio of stellar to RLOF radius (solid line), stellar to initial stellar radius (dashed line), and Roche flow overflow to initial Roche flow overflow radius (dashed dotted line).
+We note that for the majority of the evolution the ratio is larger for a three-day orbital period than a twenty-one day orbital period because of the initial orbital separation of the former is smaller than the latter in Tab.
+1.
+For simulations with an initial three day orbital period in Fig.
+2, the time of increase in stellar radius close to termination is coincident with a sharp increase in total mass loss , given in the second row.
+Furthermore, we compare the evolution of total mass loss , mass loss due to stellar wind , and mass loss due to RLOF in units of initial stellar mass .
+Before the sharp increase in stellar radius, less than Myr, the mass loss is driven by stellar winds.
+Afterwards, the mass loss is dominated by RLOF.
+This is the case for the range of stellar metallicities given in both columns.
+However, for simulations with an initial twenty-one day orbital period in Fig.
+2, we find different behavior at low and high metallicity.
+In comparing both columns of the second row, we see that at higher metallicity () RLOF is the main driver of mass loss whereas for lower metallicity () its contribution is less than the mass loss due to stellar wind.
+We note dips in the late-stage evolution of the expansion in stellar radius for simulations with significant mass loss and suspect that this may be related to the numerical challenges in modeling the stellar envelope discussed in Paxton et al.
+(2013).
+3.3 Angular momentum
+We now discuss the limit in the parameter space which gives rise to either a final increase or increase then final diminishment in the spin angular momentum of the massive star.
+For simulations with an initial three day orbital period in Fig.
+2, the third row gives the change in spin angular momentum (solid line) in units of stellar angular momentum at break-up velocity , where .
+For both metallicities, the spin-up is characterized by an increase then a sharp fall-off after the peak coincident with mass loss due to RLOF given in the second row.
+For simulations with an initial twenty-one day orbital period in Fig.
+3, this behavior is also shown in the high metallicity () case.
+The magnitude of the peak is significantly smaller in the initial orbital period of twenty-one days than three days because the tidal interaction is much weaker due to the larger orbital separation.
+Notably, we identify the limit in parameter space which gives rise to a final increase in spin angular momentum for the parameter space of low metallicity () stars interacting with a BH on an initial twenty-one day orbital period.
+In comparing the second and third row of the first column of Fig.
+3, we see an increase of without diminishment corresponding to mass loss due to RLOF that is less than mass loss due ot stellar wind.
+For initial orbital periods greater than twenty-one days, we find similar behavior with the magnitude of the peak decreasing with increasing initial orbital separation.
+The values of the final spin angular momentum of the massive star for simulations in Fig.
+1 are given in CGS units.
+The corresponding values in final specific angular momentum range from to .
 This has important implications for the spin of the BH remnant when this massive star collapses, as well as the angular momentum in the disk that forms around the BH during/just after collapse.
 We consider this in our Discussion Section below.
-3.1.2 Dependence on Mass Ratio
-We examine the spin up of the massive star and change in orbital separation over the massive star lifetime for two mass ratios , shown in Fig.
-2.
-For these simulations, we consider orbital periods of 1.5 and 3 days for a 25 massive star with a 10 and 15 BH, at stellar metallicities and .
-We find that the more extreme mass ratio encounters (with a 10 BH) lead to a larger spin up.
-This effect is more pronounced with smaller orbital periods, where stellar mass loss is greater due to the increase in encounter strength.
-This is consistent with results shown in de Mink et al.
-(2013) relating mass and angular momentum loss to mass ratio in interacting binaries.
-However, we emphasize that for all of these encounters, the massive star is tidally distorted and there is a decrease in orbital separation of the system, leading to an increase in its spin angular momentum.
-3.1.3 Dependence on Initial Spin of the Massive Star
-We consider binary encounters where the massive star is initialized over a range of rotational velocities that include non-rotating , slowly rotating , and highly rotating stars.
-We use km/s from COSMIC output, as discussed in Kenoly et al.
-(2023).
-The break-up velocity is given by km/s for , respectively.
-Fig.
-3 shows the stellar spin up and orbital separation at these varying initial rotation velocities, over the massive star lifetime.
-The evolution of and are similar for both cases of stellar metallicity.
-The spin for all three initial rotations reach close to the same value at the end of stellar lifetime.
-However, we note that when the massive star is initially rapidly rotating , this is the result of the massive star spinning down over its lifetime while increasing in orbital separation.
-3.2 Stellar Density Profiles
-In Figure 4, we show the density profile of the massive star at low and high metallicity ( and , respectively) for the case of an initial orbital period of three days.
-We expect that when tidal effects are not strong, earlier in the star’s life when the binary separation is still relatively large, the low metallicity star is more compact and the stellar density profiles are similar to single massive stars (black lines in Figure 4).
-However, when the orbital separation decreases and tidal interactions become important, the companion can have an interesting effect on the density profile of the massive star (which in turn has important implications for the density profile of the circumbinary medium).
-As we see here, tidal interactions deform the star in such a way that it loses some mass and increases in radius with the likelihood, as discussed in Cheng & Evans (2013), for shock heating in the outer layers.
-We expect that tidal interactions with an extended gas around a star may lead to ejection and therefore contribute to the density of the circumbinary environment.
-Provided that some of this gas settles into the direction into which the GRB jet (that is produced when the massive star collapses) propagates, this can lead to brighter radio emission since the radio flux scales as the square root of gas density333This is in the case of optically thin synchrotron emission.
-(Sari et al., 1998).
+Additionally, in the third row of Fig.
+2 and Fig.
+3, we show the change in spin angular momentum and the corresponding change in orbital angular momentum of the massive star in units of for each metallicity.
+In comparing these quantities, we find that for massive star – BH interactions with negligible mass loss, the decrease of matches the increase in .
+For binary interactions with significant mass loss and especially due to RLOF, we find coincident losses in the orbital angular momentum that deviate from the loss associated with the spin up of the massive star.
+3.4 Termination conditions
+We now discuss how we determine the final spin up and total mass loss of the massive star in our simulations of the interactions with a BH.
+Our conditions for termination, which set the final time for extracting these quantities, are based on tracking the formation of the helium and carbon-oxygen core as well as the central mass fraction of helium-4.
+For the sets of binary simulations corresponding to an initial three day (Fig.
+2) and twenty-one day (Fig.
+3) orbital period, we show in the fourth row the evolution of the formation of the helium () and carbon-oxygen () core in units of solar mass as well as the central mass fraction of helium-4 in units of stellar mass .
+We determine the final time of stellar termination as the coincident time where the helium an carbon-oxygen cores form and the central mass fraction of helium-4 decreases to .
+Furthermore, in the Appendix we discuss the correspondence between the formation of the carbon-oxygen core with an increase in central temperature.
+Using these conditions, present the final value of the change in spin angular momentum as well as the total mass loss in our simulations in Fig.
+1 for binary parameter space listed in Tab.
+1.
+3.5 Dependence on Metallicity
+We now discuss our results in terms of stellar metallicity in the low () and high () range of parameter space.
+In Fig.
+1, we find that within initial orbital separations (periods of 20-50 days) where significant tidal effects increase the spin angular momentum, only stars with low metallicity contribute to the parameter space of viable GRB progenitors.
+This is primarily the case because of two reasons.
+First, the stellar radius is larger for massive stars with larger metallicities and therefore they are subject to stronger tidal interactions for a given orbital period than massive stars with smaller metallicities.
+Furthermore, when the stellar radius increases at the late stages of evolution, higher metallicity massive stars will be subject to greater RLOF.
+Second, the mass loss associated with stellar winds increases with stellar metallicity.
+This combined effect leads to significant mass loss for high metallicity stars and diminished spin angular momentum below the lower limit for viable GRB progenitors.
+In Fig.
+2 and Fig.
+3, we compare results between low and high metallicity stars for the initial three day and twenty-one day orbital periods.
+In the first row, we find the stellar to RLOF radius, , is closer to unity for high metallicity massive stars throughout its lifetime.
+In the second row, we see a corresponding greater mass loss due to both stellar wind and RLOF.
+In the third row, we find that this leads to greater diminishment in spin angular momentum.
+3.6 Dependence on Mass Ratio
+We consider binary interactions between a massive star and a and BH at mass ratios, , of 1.667 and 2.5, respectively.
+We find that the range in parameter space of viable GRB progenitors is similar at both mass ratios.
+In Fig.
+1, we compare the final stellar spin up and total mass loss in binary systems with a (circle) and ("x") BH.
+We note that the spin-up of the star is smaller by a factor of for the smaller BH mass.
+The range of viable GRB candidates for low metallicity stars is within initial orbital periods of 20-40 days at both mass ratios.
+3.7 Dependence on Initial Stellar Rotation
+We consider the impact of initial rotation of the massive star in a binary with a 15 BH at a fixed initial orbital period.
+We compare simulation results initialized at three stellar velocities (, , ) under the critical limit of stellar break-up of a massive star at low () and high () metallicities.
+In Fig.
+4 and Fig.
+5, we show the massive star evolution at an initial orbital period of three and twenty-one days, respectively.
+Stellar age is given in units of years.
+In the first row, we compare the total mass loss .
+In the second row, we compare the ratio of stellar rotational velocity to initial orbital velocity .
+In the third row, we compare the change in stellar spin angular momentum with the change in orbital angular momentum in units of .
+In the fourth row, we track the evolution of the helium core mass in units of solar mass .
+For the initial orbital period of three days, in Fig.
+4, we find similar total mass loss for all three initial rotations.
+For the initial orbital period of twenty-one days, in Fig.
+5, there is greater total mass loss for the initial stellar rotation of in the low metallicity case, although at most on the order of of the initial mass of the star .
+Otherwise, the total mass loss is similar for all three initial rotations.
+For both cases in initial orbital period, we find that the stellar rotation of the massive star matches the initial rotation of the orbit at the end point of our simulations for all three initial rotations.
+In the second row of Fig.
+4 and Fig.
+5, we track the evolution of the ratio to and find that it approaches unity.
+Finally, in the third row of Fig.
+4 and Fig.
+5, we find the corresponding stellar spin angular momentum for all three initial rotations converges to a similar final spin-up value.
+This implies that in the binary parameters for stellar spin-up there is negligible dependence on initial stellar rotation because of tidal synchronization to the orbit in all cases for these initial orbital periods.
+We consider the evolution of the massive star in radial profiles of stellar density and rotation.
+In Fig.
+6, we compare profiles at early, middle, and late-stages of stellar evolution for an initially non-rotating 25 star at low metallicity () and high metallicity () at an initial orbital period of twenty-one days with a BH.
+In the top row, stellar density is given in CGS units versus stellar radius, given in solar radius .
+The bottom row shows the corresponding profile of stellar rotation in radians per second.
+At early stages, given by the solid line at stellar ages yrs, the stellar rotation profiles are close to numerical round-off error because the star is initially non-rotating.
+At middle stages, given by the dashed line at at yrs, we find that the massive star has expanded in radius close to a factor of two in comparison to early stages.
+This is consistent with the evolution of stellar radius in Fig.
+3.
+Additionally, we note that for these times in Fig.
+3 and Fig.
+5, signficant mass loss through RLOF has yet to occur.
+The stellar rotation for both low and high metallicity is higher close to the center of the star than the outer layers.
+At late stages, given by the dashed dotted line at at yrs, we find that the massive star has signficantly expanded in radius close to a factor of ten in comparison to early stages.
+In comparing this time with Fig.
+3 and Fig.
+5, we note that this stellar age is characterized by RLOF and significant mass loss in the high metallicity case only.
+The stellar rotation for the low metallicity case is uniform across the radial profile of the star.
+The stellar rotation for high metallicity case shows higher rotation at the remnant core than the outer layers.
 3.8 Additional Tests and Caveats
-There are various mechanisms, such as the Spruit-Taylor dynamo, other magnetic coupling effects, or convection effects (Spruit, 2002; Heger et al., 2005), that can transport angular momentum out of the star throughout its lifetime.
-For example, Figure 3 of Heger et al.
+There are various mechanisms, such as the Spruit-Tayler dynamo, other magnetic coupling effects, or convection effects (Spruit, 2002; Heger et al., 2005; Moreno Méndez, 2014; Ma & Fuller, 2019), that can transport angular momentum out of the star throughout its lifetime.
+For example, Fig.
+3 of Heger et al.
 (2005) shows a pronounced decrease in angular momentum when magnetic fields are present, particularly for the pre-supernova stages of a star’s evolution.
@@
 (2012) analyze the importance of the Spruit-Tayler dynamo in stars over a range of masses, and find that for stars above about 15 it becomes more difficult to sustain the dynamo, and as such the magnetic braking that normally occurs when this dynamo is active may less severe for more massive stars (particularly in the earlier stages of their evolution).
-We found that turning this flag on and off did not affect our results.
+We learned that enabling or disabling this flag, while retaining the default MESA parameters, had no impact on our results.
 We also investigated the effects of a rotating convection zone within the massive star.
 We began this investigation by setting the diffusion coefficients as prescribed in Heger et al.
-(2000) equal to 0.
+(2000) to their default values according to MESA.
 We find no change in the final angular momentum of the massive star at the end of its lifetime when this flag was turned on.
 4 Discussion
-We have investigated the properties of a massive star with a closely orbiting compact object companion and find a range of parameter space in which the massive star is spun up (i.e.
+We have investigated the properties of a massive star with a closely orbiting BH companion and find a range of parameter space in which the massive star is spun up (i.e.
 ends its life with high angular momentum).
 We conjecture that this then is a plausible model for a highly spinning BH-disk system upon the massive star’s collapse, one that is capable of producing a viable GRB.
-We note that tidal synchronization is crucial for the massive star to spin up, especially if it begins with low initial angular momentum.
 For a tidally locked binary system, the angular momentum of the massive star prior to collapse can be estimated by (Lloyd-Ronning, 2022):
@@
 The spin up of the massive star we find in our simulations in this work is consistent with this analytic result.
+We note that after collapse, the black hole central engine generally needs a spin parameter (ratio of its angular momentum to maximum possible angular momentum) of at least about 0.5 to launch a jet through the BZ process.
+For a five solar mass black hole this corresponds to roughly g cm.
+For our low metallicity simulations in particular, we find that the star at the end of its life has angular momentum in the range of to g cm.
+Our full range of binary parameter space consideration may be compared to estimates from perturbation theory which characterize the tidal encounter through a strength parameter (Press & Teukolsky, 1977), for circular orbits at stellar-BH separation distance ,
+for orbital timescale and stellar dynamical timescale .
+For a given star, encounters with smaller orbital periods will give rise to smaller or stronger tidal interactions.
+Note that the limit corresponds to the tidal disruption limit where the dynamical time scales for the orbit and star are comparable.
+In our study, we consider the weak tidal interaction regime and provide detailed simulations using the stellar evolution model described in Sec.
+2, that probe the non-linear, hydrodynamic response of a massive star throughout its lifetime.
+We find that the increase spin up of the massive star due to a decrease in orbital period are consistent with results of the spin up of polytropic stars under short time-scale interactions in parabolic orbits (Kochanek, 1992; Khokhlov et al., 1993; Cheng & Evans, 2013).
+However, we find that modeling the long-term interaction between a massive star and BH across stellar lifetime leads to the identification of a parameter region of diminished, final stellar spin angular momentum due to mass loss driven by RLOF as the stellar radius expands prior to collapse.
+These results which show a sharp decrease in stellar spin coincident with an increase in mass loss in late-stage stellar evolution.
+Paxton et al.
+(2013) discuss the challenges and uncertainty in modeling the radiation-pressure dominated envelopes in massive stars at late phases in the evolution with a 1D stellar evolution code.
+Nonetheless, we present our results at stellar termination with MESA while noting that follow-up numerical investigations that go beyond its limitations are necessary to resolve these uncertainties.
+Regarding the parameter space of encounters with competing effects that change the stellar angular momentum, Detmers et al.
+(2008) find that the tidal interaction leads to significant spin-up as well as significant losses due to mass loss due to winds and RLOF between Wolf-Rayet stars at solar metallicity and compact object companions with initial orbital periods less than 24 hours.
 We note that there is additional loss of orbital angular momentum through gravitational radiation (Peters, 1964) for which we have not accounted.
@@
 Other effects (that we do not consider in this study) such as common envelope evolution and dynamical friction can also change the angular momentum loss timescale, significantly affecting the orbital evolution and, in turn, the final angular momentum of the massive star.
-In connecting our progenitor system to a GRB - which results when the highly spinning massive star in our system collapses to a black hole - we have also not accounted for the potential loss of angular momentum upon collapse of the star.
+In connecting our progenitor system to a GRB - which results when the highly spinning massive star in our system collapses to a BH - we have also not accounted for the potential loss of angular momentum upon collapse of the star.
 Here, we are making the assumption that a highly spinning massive star at the end of its life will produce a highly spinning BH-disk system.
@@
 the BH produced from the massive star in the systems we consider) is indeed near maximum (), particularly for low metallicity systems where mass loss from line driven winds is relatively small.
+Ma & Fuller (2023) use MESA to investigate the interaction between a WR star with a BH companion.
+They find that a WR star at solar metallicity experiences significant angular momentum loss due to stellar winds.
+In contrast, WR stars with low metallicity retain their angular momentum and are spun up due to tidal interactions.
 Therefore, under the reasonable assumption of a highly spinning massive star produces a highly spinning BH-disk system, this central engine with high angular momentum will then produce a longer lasting GRB jet.
@@
 We suggest that in some cases (especially for higher metallicity stars with a “puffier” density profile or for more vigorous tidal interactions), a denser (relative to a single massive star) medium may be formed around the binary system to a larger radius and may potentially lead to a brighter radio afterglow, where the radio flux scales as roughly the square root of the circumbinary density (Sari et al., 1998).
-The tidal interaction between the black hole and the massive star is the underlying mechanism of the spin up of our massive star, as well as the extended density profile that deviates from the typical profile of a single massive star.
+The specifics of the mass-loss from line driven winds in these systems will also affect the circumbinary density profile.
+In this paper, we simply offer the suggestion of a varying circumbinary environment depending on the resulting density profiles of star, as seen in Figure 6, but leave a detailed investigation of this issue to a following publication.
+The tidal interaction between the BH and the massive star is the underlying mechanism of the spin up of our massive star, as well as the extended density profile that deviates from the typical profile of a single massive star.
 We note that Sciarini et al.
@@
 5 Conclusions
-We have modeled the evolution of a massive star with a closely orbiting compact object companion.
+In this paper we have explored the influence of a BH companion on a massive star throughout its lifetime, over a range of parameter space varying orbital period, metallicity, initial spin of the star, mass ratio, accretion rate, magnetic fields, and numerical stopping conditions.
 Our study is motivated by the interesting result that radio bright GRBs tend to have a longer prompt gamma-ray burst duration compared to radio dark GRBs.
 As such, we have explored systems that potentially lead to a massive star with more angular momentum (and therefore a longer lived accretion disk and jet which reflects the duration of the prompt emission).
-We have also discussed the dependence of the density profile of the massive star on binary interaction and stellar metallicity (which has implications for the density of the circumbinary medium).
+We have focused on a range of parameter space in which the initial orbital period (binary separation) is such that the massive star does not initially experience RLOF and tidal interactions still play a significant role in the spin-angular momentum evolution of the massive star.
+We have also briefly discussed the dependence of the density profile of the massive star on binary interaction and stellar metallicity (which has implications for the density of the circumbinary medium).
 Our main conclusions are as follows:
-We find that massive stars with closely orbiting compact object companions are significantly spun up through tidal interactions, and end their life with high angular momentum, conducive to producing a highly spinning BH-disk system that could power a relatively long-lived GRB jet.
+We identify a particular range of parameter space where massive stars with closely orbiting BH companions are significantly spun up through tidal interactions, and end their life with high angular momentum, conducive to producing a highly spinning BH-disk system that could power a relatively long-lived GRB jet.
 The specific angular momentum of the star at the end of its life aligns with the expectation of the angular momentum needed for a roughly 10 star to launch a relativistic jet through the Blanford-Znajek process (Blandford & Znajek, 1977).
-We find tidal effects leading to significant stellar spin up for orbital periods up to about 4 days, and over a range of mass ratios.
-We find that the tidal spin up is more pronounced for higher metallicity stars for a given binary orbital period, due to the larger radius/extent of the star.
-However, the spin up is still significantly present in low metallicity systems.
-We find that spin up depends strongly on the initial rotation of the massive star.
-Unsurprisingly, stars with more initial angular momentum are not spun up for the orbital periods we consider.
-In these cases, tidal syncing causes the massive star to spin down to synchronize with the orbital period of the system.
-We present two sample density profiles of a massive star at two metallicities, illustrating the greater extent of the higher metallicity star.
-We discuss how this not only affects tidal interactions but also potentially impacts the density profile of the medium surrounding the binary.
+We identify low metallicity () massive stars interacting with both 15 and 10 BH for an initial orbital period of 20-50 days as viable progenitor binary systems.
+For this range of parameter space, we note significant final stellar spin-up and negligible final total mass loss on the order of a factor of of the initial mass of the star.
+Outside of this region, non-viable binaries are characterized by significant final mass loss close to a factor of unity or at an orbital separation where the tidal interaction is negligible ( 60).
+Additionally, we note a very slight decrease in orbital separation on the order of for simulations with negligible mass loss including the viable progenitor binary systems.
+We find that the final stellar spin angular momentum does not depend on the initial rotation of the massive star for the parameters we have investigated.
+In these cases, the tidal effects will spin up or spin down the massive star in order to synchronize with the orbital period of the system.
 The results of our work here inform future detailed studies into the physics of gamma-ray bursts.
@@
 Acknowledgments
+We thank the referee for the detailed and constructive feedback.
 We thank the MESA community for helpful conversations and the many resources provided to help run MESA.